OpenStudy (anonymous):

Can somebody teach me how do find the quotient in this: -6y^3+84y^2+27y and its all over -3y

OpenStudy (kirbykirby):

You can use the rule: \[\frac{x+y}{a}=\frac{x}{a}+\frac{y}{a}\] You can also use power rules: \[\frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (kirbykirby):

So, \[\frac{-6y^3+84y^2+27y}{-3y}\] Can you break this up into 3 fractions as mentioned in the first rule?

OpenStudy (anonymous):

Thanks, anus head